# Deep Reinforcement Learning for Image-to-Image Translation

### 14 Dec 2023 v2

@@ Abstract
Image to Image Translation, Deep Reinforcement Learning, Meta Policy, Auxiliary Learning.
@@ 3.2 Stochastic Meta Policy of Planning and Acting
Our RL-I2IT framework is designed to handle high-dimensional continuous states and actions in an infinite-horizon Markov Decision Process (MDP). It incorporates a novel component, a planner, specifically for continuous plan space. This approach diverges from traditional policies that directly map environmental states to actions [27]. Instead, it bifurcates the mapping process into two distinct steps: first state $\rightarrow$ plan, and then plan $\rightarrow$ action. We term this new two-step mapping process as a “meta policy”, which allows for more intricate and layered decision-making compared to standard reinforcement learning models. We define the new mapping process as meta policy, and the new MDP for our RL-I2IT can be represented by the tuple $({\cal S},{\cal P},{\cal A},{\cal U},{r},{\gamma})$. ${\cal S}$ is a set of states, ${\cal P}$ is a continuous plan, ${\cal A}$ is a continuous action, and ${\cal U}:{\cal S}\times{\cal P}\times{\cal S}\times{\cal A}\rightarrow[0,\infty)$ represents the state transition probability density of the next state ${\bf s}_{t+1}$ given state ${\bf s}_{t}\in{\cal S}$, plan ${\bf p}_{t}\in{\cal P}$ and action ${\bf a}_{t}\in{\cal A}$.
@@
Formally, suppose a meta policy is defined as ($\kappa,\pi$). The stochastic plan is modeled as a subspace of the deformation field that gives a low-dimensional vector ${\bf p}_{t}$ based on the state ${\bf s}_{t}$, while action ${\bf a}_{t}$ is determined by the plan ${\bf p}_{t}$. Consider a parameterized planner $\kappa_{\psi}$ and actor $\pi_{\phi}$, the stochastic plan is sampled as a representation: ${\bf p}_{t}\sim\kappa_{\psi}({\bf p}_{t}|{\bf s}_{t})$, and the action is generated by decoding the plan vector ${\bf p}_{t}$ into a high-dimensional executable action: ${\bf a}_{t}=\pi_{\phi}({\bf a}_{t}|{\bf p}_{t})$. In practice, we reparameterize the planner and stochastic plan jointly using a neural network approximation ${\bf p}_{t}=f_{\psi}({\bf\epsilon}_{t},{\bf s}_{t})$, known as the reparameterization trick [2], where ${\bf\epsilon}_{t}$ is an input noise vector sampled from a fixed Gaussian distribution. Moreover, we maximize the entropy of the plan to improve exploration and robustness. The augmented objective function is formulated as follows:
@@ 3.3 Learning Planner and Critic
The derivation for the case of the critic evaluating the actor can be found in Appendix A. Our experimental results to be reported in Table X will show that the critic evaluates actor’s action results in an inferior performance to that the critic evaluates the planner.
@@ 3.5 Environment Settings in Practice
In the given RL-I2IT framework, environment designs are tailored for various applications, with detailed guidance in each section. This section outlines general principles for selecting rewards, focusing on the critic’s role in evaluating plans rather than actions. The plans, being a subset of potential actions, serve as high-level instructions for the actor to create specific actions. Evaluation measures for structural or global image information, such as SSIM or the DICE score [29], are proposed as rewards for assessing these plans, as the evaluations shown in Table IX. However, the text emphasizes that the choice of reward should remain flexible and be empirically tested in the context of individual applications.
@@ 4.1.1 RL-I2IT Setting
For digits transform, the state is a concatenation of the predicted image and the target image. The Dice score [29] is used as the reward, and the NCC loss [31, 32] is leveraged for auxiliary learning. A very simple network structure is used to construct the planer, the actor, and the critic, as shown in Fig. 3. More specifically, the plan is a one-channel $7\times 7$ feature map (49-dimensional plan), and the actor outputs a deformation field, which is used to transform the source image by spatial transformer network (STN) [33]. All convolution operations use a $3\times 3$ kernel with the LeakyReLU activation function. The downsampling operation is performed by max-pooling, and all upsampling operations are performed with the nearest interpolation.
@@ 4.2.1 RL-I2IT Setting
Figure 8: Visual comparison of our RL-I2IT with pix2pix, PAN, pix2pixHD, and DRPAN over all tasks.
@@ 4.4 Image Style Transfer
Our RL-I2IT framework provides a good solution for NST. It can be used to learn a lightweight NST model that is applied iteratively for NST. To preserve spatial structures of images, the latent plans in our model are sampled from a 2D Gaussian distribution that is estimated by the actor and forwarded to the executor to generate intermediate images. In addition, we develop a Fully Convolutional Network (FCN) based planner-actor structure so that the model can process input images of any size. Fig. 10 shows some examples of our step-wise NST. We can see that our RL-based step-wise method tends to preserve more details and structures of the content image in early steps and synthesize more style patterns in later steps, resulting in a more flexible control of the stylization degree. Furthermore, our model is a lightweight and flexible NST model compared to existing methods, making it more efficient computationally. To the best of our knowledge, this is the first work that successfully leverages RL for the NST scenario.
@@ 4.4.1 RL-I2IT Setting
Style Learning. To make the moving image not deviate from the content image, the model trains the planner and the actor based on collected training data from the agent-environment interaction and changes dynamically in experience replay. More specifically, the planner and actor form a conditional generative process that translates state ${\bf s}_{t}$ to output moving image ${\bf m}_{t}$ at time $t$. Note that ${\bf s}_{t}$ is initialized to content image ${\bf c}$ and ${\bf s}_{t+1}$ is equivalently ${\bf m}_{t}$. Inspired by [53], we apply the content loss $\mathcal{L}^{CO}$, style loss $\mathcal{L}^{ST}$, and total variation regularization $\mathcal{L}^{TV}$ to optimize the model parameters of planner and actor. These losses can better measure perceptual and semantic differences between the moving image and content image ${\bf c}$.
Content Loss $\mathcal{L}^{CO}$. Following [53], we use a pre-trained neural network $F$ to extract the high-level feature representatives of ${\bf m}_{t}$ and ${\bf c}$. The reason for using this $F$ is to encourage moving image ${\bf m}_{t}$ to be perceptually similar to content image ${\bf c}$ but does not force them to match exactly. Denote $F^{j}(\cdot)$ as the activations of the $j$-th layer of $F$. Suppose $j$-th layer is a convolutional layer, then the output of $F^{j}(\cdot)$ will be a feature map with size $C^{j}\times H^{j}\times W^{j}$, where $C^{j}$, $H^{j}$, and $W^{j}$ represent the number of channels, height, and width in the feature map of layer $j$, respectively. We apply the Euclidean distance, which is squared and normalized, to design the content loss as follows,
@@
Style Loss $\mathcal{L}^{ST}$. To penalize ${\bf m}_{t}$ when it deviates in content from ${\bf c}$ and in style from ${\bf e}$, we define by following [54] a Gram matrix $G^{j}({\bf x})=\frac{\tilde{F}^{j}({\bf x})(\tilde{F}^{j}({\bf x}))^{\top}}{C^{j}H^{j}W^{j}}\in\mathbb{R}^{C^{j}\times C^{j}}$, where $\tilde{F}^{j}(\cdot)$ is obtained by reshaping $F^{j}(\cdot)$ into the shape $C^{j}\times H^{j}W^{j}$. The style loss can be defined as a squared Frobenius norm of the difference between the Gram matrices of ${\bf m}_{t}$ and ${\bf e}$. To preserve the spatial structure of images, we use a set of layers, $J$, instead of a single layer $j$. Thus, we define the style loss to be the sum of losses for each layer $j\in J$ ($|J|=4$ in our experiments):
@@
Total Variation Regularization $\mathcal{L}^{TV}$. To ensure spatial smoothness in moving image ${\bf m}$, we use a total variation regularizer $\mathcal{L}^{TV}({\bf m}_{t})$, which has been widely used [55, 53].
Putting all components together, the final style loss is
@@ 4.4.2 Experiment
Ablation Study. (1) We study the effect of the RL model in our framework. As shown in Fig. 14, compared with the method that only uses Planner-Actor (PA), our method can generate more stable and clear stylized images at step 1. At step 10, PA loses the content information completely without the help of RL (ours w/o RL), while our method can still produce amazing results. We also show the corresponding quantitative comparison in Table VI. We can easily see that at both steps 1 and 10, our model achieves the best performance consistently. This indicates RL can improve the performance of DL-based NST models. (2) Since AdaIN and StyTR2 methods can adjust the hyperparameter ‘alpha’$\in[0,1]$ and the round of repetitive stylization to control the degree of stylization in the final results, respectively, we compare our method with them accordingly in Fig. 15. From the visualization results, we can see that the results of AdaIN are in the under-stylized state even if the style control hyperparameter is changed. Moreover, StyTR2 gets the results with small style changes and low quality after multiple rounds. However, our method not only ensures the gradual change in style, but also produces very smooth results. (3) We evaluate the necessity of the high-dimension for style transfer, showing in Fig. 16. From the result, it can be seen that 1D latent vector can only generate a single style pattern and cannot retain the semantic information of the content. In contrast, high-dimensional latent tensor preserves the structure and semantic information of content, allowing for simultaneous stylization and content reconstruction.
@@
Figure 18: Illustration of our step-wise video style transfer process using the RL-I2IT framework. The frames are stylized stronger with the perdition steps smoothly. The model tends to preserve more details and structures of the content in the early steps and synthesize more style patterns in the later steps. Our model is capable of generating stable stylized results across different degrees of stylization.
@@ 4.5 Video Style Transfer
The recent works in the field of video style transfer [65, 66, 67, 64, 64] have been successful. However, these methods are facing the challenge of limited diversity. They can only generate results with a singular degree of stylization, without considering the creation of stylized videos tailored to different audiences with varying degrees of stylization.
To achieve a diverse range of stylization levels in video style transfer, we employ the RL-I2IT framework. Specifically, we have made adjustments to the neural network for image style transfer. By using a CNN+RNN architecture[68] for the planner and actor for both frame-wise and step-wise smoothing, our model can perform video NST tasks. Fig. 18 illustrates some examples of our step-wise video Neural Style Transfer. We observe that our RL-I2IT framework approach not only retains the advantages seen in image style transfer, namely the tendency to preserve more details and structure of the frame in the early steps while synthesizing more style patterns in the later steps but also generates stable outputs across different degrees of stylization.
@@ 4.5.1 RL-I2IT Setting
In video style transfer, we maintain most of the settings used in image style transfer. The difference lies in initializing the moving images using frames.
Building upon the image style transfer network, we made slight modifications. The network’s details are illustrated in Fig. 19. We introduced two GRUs: the step-wise GRU and the frame-wise GRU. The step-wise GRU retains information between steps, ensuring smoother step-wise stylization. Meanwhile, the frame-wise GRU preserves information between frames, enforcing consistency in style patterns among adjacent frames.
Compound Temporal Regularization $\mathcal{L}^{CT}$. Inspired by [67], we add a compound temporal regularization for video style transfer. Specifically, we first generate motions $M(\cdot)$ and then synthesize adjacent frames. With this approach, we do not need to estimate optical flow in the training process and we can guarantee the optical flows are absolutely accurate. Given noise $\triangle$, to maintain temporal consistency, we can minimize the following loss
@@ 4.5.2 Experiment
(2) Similarly, we have compared our method with the results produced when Step-wise GRU is removed, and the results are shown in Fig. 21 (second and third rows). We can clearly see that most of the face details of the protagonist and dragon have been lost at step 10 when using our method without the Step-wise GRU. Also, the external details of the protagonist and dragon are completely lost in step 20. Our method with using Step-wise GRU, on the other hand, obtains very fine results even at step 20. (3) Table VIII shows the comparison of the temporal loss of our method with Frame-wise GRU (FWG) and without FWG. We find that the temporal loss is very low if we use FWG, which means the obtained final results are more consistent from frame to frame. The above experiments have shown that RL, Step-wise , and Frame-wise GRU all greatly improve the performance of the model.
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grants 61602065 and 42130608, Sichuan province Key Technology Research and Development project under Grant 2021YFG0038.

### 24 Feb 2024 v4

# Image-to-Image Translation with Deep Reinforcement Learning

@@ Abstract
Image to Image Translation, Deep Reinforcement Learning, Meta Policy, Deep Learning, Generative Model.
@@ 3.2 Stochastic Meta Policy of Planning and Acting
Our RL-I2IT framework is designed to handle high-dimensional continuous states and actions in an infinite-horizon Markov Decision Process (MDP). It incorporates a novel component, a planner, specifically for continuous plan space. This approach diverges from traditional policies that directly map environmental states to actions [27]. Instead, it bifurcates the mapping process into two distinct steps: first state $\rightarrow$ plan, and then plan $\rightarrow$ action. We term this new two-step mapping process as a “meta policy”, which allows for more intricate and layered decision-making compared to standard reinforcement learning models. The new MDP for our RL-I2IT can be represented by the tuple $({\cal S},{\cal P},{\cal A},{\cal U},{r},{\gamma})$. ${\cal S}$ is a set of states, ${\cal P}$ is a continuous plan, ${\cal A}$ is a continuous action, and ${\cal U}:{\cal S}\times{\cal P}\times{\cal S}\times{\cal A}\rightarrow[0,\infty)$ represents the state transition probability density of the next state ${\bf s}_{t+1}$ given state ${\bf s}_{t}\in{\cal S}$, plan ${\bf p}_{t}\in{\cal P}$ and action ${\bf a}_{t}\in{\cal A}$.
@@
Formally supposing a meta policy is defined as ($\kappa,\pi$). The stochastic plan is modeled as a subspace of the deformation field that gives a low-dimensional vector ${\bf p}_{t}$ based on the state ${\bf s}_{t}$, while action ${\bf a}_{t}$ is determined by the plan ${\bf p}_{t}$. Consider a parameterized planner $\kappa_{\psi}$ and actor $\pi_{\phi}$, the stochastic plan is sampled as a representation: ${\bf p}_{t}\sim\kappa_{\psi}({\bf p}_{t}|{\bf s}_{t})$, and the action is generated by decoding the plan vector ${\bf p}_{t}$ into a high-dimensional executable action: ${\bf a}_{t}=\pi_{\phi}({\bf a}_{t}|{\bf p}_{t})$. In practice, we reparameterize the planner and stochastic plan jointly using a neural network approximation ${\bf p}_{t}=f_{\psi}({\bf\epsilon}_{t},{\bf s}_{t})$, known as the reparameterization trick [2], where ${\bf\epsilon}_{t}$ is an input noise vector sampled from a fixed Gaussian distribution. Moreover, we maximize the entropy of the plan to improve exploration and robustness. The augmented objective function is formulated as follows:
@@ 3.3 Learning Planner and Critic
The derivation for the case of the critic evaluating the actor can be found in Appendix A. Besides, our experimental results (as reported in Table X) also show that the critic evaluates actor’s action results in an inferior performance to that the critic evaluates the planner.
@@ 3.5 Environment Settings in Practice
In the given RL-I2IT framework, environment designs are tailored for various applications, with detailed guidance in each section. This section outlines general principles for selecting rewards, focusing on the critic’s role in evaluating plans rather than actions. The plans, being a subset of potential actions, serve as high-level instructions for the actor to create specific actions. Evaluation measures for structural or global image information, such as SSIM [29] or the DICE score [30], are proposed as rewards for assessing these plans. However, the choice of reward remains flexible and should be empirically tested in the context of individual applications.
@@ 4.1.1 RL-I2IT Setting
For digits transform, the state is a concatenation of the predicted image and the target image. The Dice score is used as the reward, and the NCC loss is leveraged for auxiliary learning [32, 33]. A very simple network structure is used to construct the planer, the actor, and the critic, as shown in Fig. 3. More specifically, the plan is a one-channel $7\times 7$ feature map (49-dimensional plan), and the actor outputs a deformation field, which is used to transform the source image by spatial transformer network (STN) [34]. All convolution operations use a $3\times 3$ kernel with the LeakyReLU activation function. The downsampling operation is performed by max-pooling, and all upsampling operations are performed with the nearest interpolation.
@@ 4.2.1 RL-I2IT Setting
Figure 8: Visual comparison of our RL-I2IT with pix2pix, PAN, pix2pixHD, and DRPAN over photo translation tasks.
@@ 4.4 Image Style Transfer
Our RL-I2IT framework provides a good solution for NST. It can be used to learn a lightweight NST model that is applied iteratively for NST. To preserve spatial structures of images, the latent plans in our model are sampled from a 2D Gaussian distribution that is estimated by the planner and forwarded to the actor to generate intermediate images. In addition, we develop a Fully Convolutional Network (FCN) based planner-actor structure so that the model can process input images of any size. Fig. 10 shows some examples of our step-wise NST. We can see that our RL-based step-wise method tends to preserve more details and structures of the content image in early steps and synthesize more style patterns in later steps, resulting in a more flexible control of the stylization degree. Furthermore, our model is a lightweight and flexible NST model compared to existing methods, making it more efficient computationally. To the best of our knowledge, this is the first work that successfully leverages RL for the NST scenario.
@@ 4.4.1 RL-I2IT Setting
Style Learning. To make the moving image not deviate from the content image, the model trains the planner and the actor based on collected training data from the agent-environment interaction and changes dynamically in experience replay. More specifically, the planner and actor form a conditional generative process that translates state ${\bf s}_{t}$ to output moving image ${\bf m}_{t}$ at time $t$. Note that ${\bf s}_{t}$ is initialized to content image ${\bf c}$ and ${\bf s}_{t+1}$ is equivalently ${\bf m}_{t}$. Inspired by [54], we apply the content loss $\mathcal{L}^{CO}$, style loss $\mathcal{L}^{ST}$, and total variation regularization $\mathcal{L}^{TV}$ to optimize the model parameters of planner and actor. These losses can better measure perceptual and semantic differences between the moving image and content image.
Content Loss. Following [54], we use a pre-trained neural network $F$ to extract the high-level feature representatives of ${\bf m}_{t}$ and ${\bf c}$. The reason for using this $F$ is to encourage moving image ${\bf m}_{t}$ to be perceptually similar to content image ${\bf c}$ but does not force them to match exactly. Denote $F^{j}(\cdot)$ as the activations of the $j$-th layer of $F$. Suppose $j$-th layer is a convolutional layer, then the output of $F^{j}(\cdot)$ will be a feature map with size $C^{j}\times H^{j}\times W^{j}$, where $C^{j}$, $H^{j}$, and $W^{j}$ represent the number of channels, height, and width in the feature map of layer $j$, respectively. We apply the Euclidean distance, which is squared and normalized, to design the content loss as follows,
@@
Style Loss. To impose penalties on ${\bf m}_{t}$ when it deviates in content from ${\bf c}$ and in style from ${\bf e}$, we define the style loss by following [55] a Gram matrix $G^{j}({\bf x})=\frac{\tilde{F}^{j}({\bf x})(\tilde{F}^{j}({\bf x}))^{\top}}{C^{j}H^{j}W^{j}}\in\mathbb{R}^{C^{j}\times C^{j}}$, where $\tilde{F}^{j}(\cdot)$ is obtained by reshaping $F^{j}(\cdot)$ into the shape $C^{j}\times H^{j}W^{j}$. The style loss can be defined as a squared Frobenius norm of the difference between the Gram matrices of ${\bf m}_{t}$ and ${\bf e}$. To preserve the spatial structure of images, we use a set of layers, $J$, instead of a single layer $j$. Thus, we define the style loss to be the sum of losses for each layer $j\in J$ ($|J|=4$ in our experiments):
@@
Total Variation Regularization. To ensure spatial smoothness in moving image ${\bf m}_{t}$, we use a total variation regularizer $\mathcal{L}^{TV}$.
Putting all components together, the final style learning loss is
@@ 4.4.2 Experiment
Ablation Study. (1) We study the effect of the RL model in our framework. As shown in Fig. 14, compared with the method that only uses Planner-Actor (PA), our method can generate more stable and clear stylized images at step 1. At step 10, PA loses the content information completely without the help of RL (ours w/o RL), while our method can still produce amazing results. We also show the corresponding quantitative comparison in Table VI. We can easily see that at both steps 1 and 10, our model achieves the best performance consistently. This indicates RL can improve the performance of DL-based NST models.
(2) Since AdaIN and StyTR2 methods can adjust the hyperparameter ‘$\alpha$$\in[0,1]$ and the round of repetitive stylization to control the degree of stylization in the final results, respectively, we compare our method with them accordingly in Fig. 15. From the visualization results, we can see that the results of AdaIN are in the under-stylized state even if the style control hyperparameter is changed. Moreover, StyTR2 gets the results with small style changes and low quality after multiple rounds. However, our method not only ensures the gradual change in style, but also produces very smooth results.
(3) We evaluate the necessity of the high-dimension for style transfer, showing in Fig. 16. From the result, it can be seen that 1D latent vector can only generate a single style pattern and cannot retain the semantic information of the content. In contrast, high-dimensional latent tensor preserves the structure and semantic information of content, allowing for simultaneous stylization and content reconstruction.
@@
Figure 18: Illustration of our step-wise video style transfer process using the RL-I2IT framework. The first row and the second row represent two different frames from the same video sequence, respectively. The frames are stylized stronger with the perdition steps smoothly. The model tends to preserve more details and structures of the content in the early steps and synthesize more style patterns in the later steps. Our model is capable of generating stable stylized results across different degrees of stylization. Meanwhile, at the same level of stylization, the same object exhibits similar stylized characteristics in two different frames.
@@ 4.5 Video Style Transfer
The recent works in the field of video style transfer [65, 66, 67, 64] have been successful. However, these methods are facing the challenge of limited diversity. They can only generate results with a singular degree of stylization, without considering the creation of stylized videos tailored to different audiences with varying degrees of stylization.
To achieve a diverse range of stylization levels in video style transfer, we employ the RL-I2IT framework. Specifically, we have made adjustments to the neural network for video style transfer. By using a CNN+RNN architecture[68] for the planner and actor for both frame-wise and step-wise smoothing, our model can perform video NST tasks. Fig. 18 illustrates some examples of our step-wise video NST. We observe that our RL-I2IT framework approach not only retains the advantages seen in image style transfer, namely the tendency to preserve more details and structure of the frame in the early steps while synthesizing more style patterns in the later steps, but also generates stable outputs across different degrees of stylization.
@@ 4.5.1 RL-I2IT Setting
In video style transfer, we maintain most of the settings used in image style transfer. The difference lies in initializing the moving images using frames. In addition, building upon the image style transfer network, we made slight modifications for video style transfer. Two GRUs are introduced: the step-wise GRU and the frame-wise GRU. The step-wise GRU retains information between steps, ensuring smoother step-wise stylization. Meanwhile, the frame-wise GRU preserves information between frames, enforcing consistency in style patterns among adjacent frames. The network’s details are illustrated in Fig. 19.
Compound Temporal Regularization. Inspired by [67], we add a compound temporal regularization for video style transfer. Specifically, we first generate motions $M(\cdot)$ and then synthesize adjacent frames. This innovative approach eliminates the necessity to estimate optical flow in the training process and it can be guaranteed that the optical flows are absolutely accurate. We also include random noise $\triangle$ to maintain temporal consistency, and the compound temporal regularization is defined as:
@@ 4.5.2 Experiment
(2) We compared our method with the results produced when Step-wise GRU is removed, and the results are shown in Fig. 21 (second and third rows). We can clearly see that most of the face details of the protagonist and dragon have been lost at step 10 when using our method without the Step-wise GRU. Also, the external details of the protagonist and dragon are completely lost in step 20. Our method with using Step-wise GRU, on the other hand, obtains very fine results even at step 20.
(3) Table VIII shows the comparison of the temporal loss of our method with Frame-wise GRU (FWG) and without FWG. We find that the temporal loss is very low if we use FWG, which means the obtained final results are more consistent from frame to frame. The above experiments have shown that RL, Step-wise and FWG all greatly improve the performance of the model.
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grants 42375148 and 42130608, Sichuan province Key Technology Research and Development project under Grants 2023YFG0305.